J. \underline{List of Preprocessor Options}

Numerical Options The user must select the first option and any or all of the remaining options. congrad5pt - Use 5-point conjugate-gradient elliptic solver for barotropic streamfunction. islands - Include islands in relaxation topography. fourfil - Fourier filter the prognostic variables in the longitudinal direction at high latitudes. northpole - No island at the North Pole; see Section H. taperhrzvd - Taper horizontal and isopycnal mixing coefficients at high latitudes; see Section H. upwind3 - Use third-order upwinding option; see Section H.

Model Resolution Default model resolution is 2×2. Option model3×3 has been extensively tested; option model1×1 has not. model1×1 - Select default parameter values for 1×1 resolution. model3×3 - Select default parameter values for 3×3 resolution.

Vertical Sub-grid-scale Mixing Schemes One of the first two options must be selected. If kmix is selected, then implicitvmix must also be selected. Multiple sub-options, which are indented following the major option, may be selected. constvmix - Constant vertical profiles for fkpm and fkph. implicitvmix - Implicit vertical mixing; convective adjustment is bypassed. kmix - Mixing scheme based on the K-profile planetary boundary layer scheme by Large, McWilliams, and Doney; see Section G.

kmixcheckekmo - Check Ekman Layer and Monin-Obukhov length scales.

kmixnori - The vertical viscosity and diffusivity below the boundary layer do not depend on the local Richardson number, but are set to the background values, fkpm and fkph. kmixdd - Include double-diffusion contribution to vertical diffusivity.

J-1 Horizontal and Isopycnal Sub-grid-scale Mixing Schemes One of the first two options must be selected. In addition, isopycmix may be selected. biharmonic - Constant coefficient biharmonic horizontal mixing (∇^{4}). consthmix - Constant coefficient Laplacian horizontal mixing (∇^{2}). isopycmix - Mesoscale eddy parameterization; see Section F. This only applies to the tracer equations, and it is recommended that the coefficient ah is zero when this option is chosen.

isopycmixspatialvar - Taper the isopycnal mixing coefficient to zero when isopycnal slopes become steep. This option is valid only when isopycmix is set.

Surface Forcing Options The surface forcing options are described in Section I. Only one of the first four major options should be selected. sflxiceflx and sflxsw may be selected in addition. sflxcombo - Specify type of tracer forcing data: This is a combination of air-sea fluxes (based on air temperature, specific humidity, solar short wave flux, cloud fraction, precipitation, and windspeed) and restoring terms (due to restoring fields of SST and SSS).

sflxcomboavan - Compute annual averages of several surface tracer fluxes for diagnostic purposes. restormed - Restore sea-surface temperature and salinity back to specified values in the Mediterranean Sea (the 3×3 model does not have a Mediterranean Sea).

sflxpvm - Provide all surface fluxes via the Flux Coupler.

sflxpvmicetilt - Compute the tilt of the sea surface; necessary for some ice models. restorst - Restore sea-surface temperature and salinity back to specified values globally. sflxdataflx - Specify all fluxes from data: forcing data consist of non-solar heat flux, solar heat flux, and net freshwater flux. sflxiceflx - Check for ice formation in the uppermost kmxice layers in the ocean model, change temperature and salinity if ice forms or melts, and keep track of previously formed ice. Can be used either for ocean only model runs and should be used for coupled runs. sflxsw - Treat solar short wave flux differently from all other surface heat fluxes: this flux appears as a right-hand-side term in sourct and is allowed to penetrate the water with an exponentially decreasing absorption factor. The Jerlov water type, jwatertype, is specified in blkdta.F.

Boundary Conditions cyclic - Impose cyclic longitudinal boundary conditions.

J-2 Input/Output crayio - Use I/O based on Cray getwa and putwa commands for out-of-core computations. diskless - Simulate disk storage using memory array for in-core computations.

Model Performance and Monitoring multitasking - Allow parallel processing via latitude bands.

readchunk - Specify the chunk size for each set of latitudes so that the computational load is balanced among all of the processors. Presently, this option is valid only when using the 2×2 model, with ntasks=4,8, or 16; the chunks are defined in blkdta.F.

keepterms - Use arrays to retain terms in momentum and tracer equations, instead of recomputing via statement functions. testcfl - Test whether any velocity exceeds the local CFL criterion by a factor of cflcrt. timing - Compute the time spent per gridpoint per timestep; report cpu and wall-clock times. Valid for Cray Y-MP systems.

Conveniences autorestart - Provide an automatic restart from the latest model restart files. readrmsk - Read in region information from the file “regionmask”. time average - Create time averages of the model fields, forcing and poleward transports, see Section H.

Hardware Options NOMSS - Do not use the NCAR Mass Storage System subroutines msread and mswrite to move history and restart files to archival storage; rather, use replacement subroutines which use Unix commands via “call system” or “ishell” to move these files to a more permanent location. CRAY - Indicates the ocean model is being run on a Cray computer; “ishell” is used for all system calls from the model and Cray pointers are used in the subroutine step.

Internal Options hmixalreadyset - An internal option set in the files fdifm.h and fdift.h ioalreadyset - An internal option set in the file odam.F

J-3 Unsupported Options The following options exist in the NCAR CSM ocean model, but they are not supported by the NCAR CSM developers and should be activated by the user only after careful testing to ensure correctness. congrad9pt - Use 9-point conjugate-gradient elliptic solver for barotropic streamfunction. constvmixdmxl - Enhance vertical diffusivity in the top 50m (3 or 4 layers). constvmixvmxl - Enhance vertical viscosity in the top 50m (3 or 4 layers). coupled ocean only - Run the equivalent of the stand-alone ocean model in coupled mode, through the flux-coupler. This option requires the existence of special climatological ice, land, and atmospheric component models, which are presently unavailable. coupled ocean ice - Run the ocean model coupled with an active ice model through the flux-coupler. This option requires special climatological land and atmospheric component models, which are presently unavailable. extras - Provide extra diagnostic information. fio - Fortran direct-access I/O. firfil - Finite-impulse filter. insitu - Compute polynomial expansion coefficients for the equation of state using in-situ temperature. knudsen - Use Knudsen equation of state for sea water. multitaskqueue - Monitor the status of the NCAR Cray multitasking queue; shut down and exit upon receiving the proper signal. nlhmix - Nonlinear horizontal sub-grid-scale mixing. nohilats - Drop metric nonlinear terms in the momentum equation. ppvmix - Use Pacanowski and Philander vertical mixing. sflxtharm - Use annual and first two harmonics of the seasonal cycle for restoring of the surface tracer fields. sflxwharm - Use annual and first two harmonics of the seasonal cycle for wind stress forcing. skipland - Perform computations only over ocean points. symmetry - Symmetric boundary condition about the northern model boundary. tcvmix - Mellor-Yamada turbulence closure scheme.

lalg - tcvmix suboption. leq - tcvmix suboption.

J-4